# iIntel Labs, USA

### v1

@@ 4.2 Quantization
It has been well established by prior work Dong et al. (2023); Jiang et al. (2025); Yang et al. (2024) that outlier channels in ViT activations follow a fixed pattern, allowing anticipation in advance during calibration. However, unlike ViTs, VMMs exhibit dynamic variations in activation outlier channels across time-steps. As shown in Figure 3, a comparison of activations $\bar{A},\bar{B}$ at time-steps 0 and 130 reveals that outlier channels (in red) differ significantly between time-steps. This temporal variance in outlier positions suggests that a static, pre-determined calibration approach, as used in ViTs, is insufficient for VMM quantization. Therefore, VMM necessitates dynamic identification of outlier channels at inference to reduce quantization error.
@@ 5 OuroMamba DFQ Framework
Input: A pre-trained FP VMM model $P$ with $L$ layers, Gaussian noise batch $X_{\mathcal{B}}$, task-specific targets $T_{G_{\mathcal{B}}}$, neighborhood size $\mathcal{N}$, iterations $G$.
Output: A set of generated synthetic samples $X^{*}_{\mathcal{B}}$.
for $g=1,2,\dots,G$ do
       Input $X_{\mathcal{B}}$ into $P$ ;
       for $l=1,2,\dots,L$ do
             Capture per-time-step original $h^{l}(t)$, $\Delta^{l}(t)$ ;
             Set $w^{l}_{t}=\texttt{mean}_{E}(\Delta^{l}(t))$ ;
             Compute $h_{p}^{l}(t)$;
             Extract implicit attention;
             Compute $\mathcal{L}^{C}_{l}=\sum_{t}\mathcal{L}^{C}_{l,t}$;
       end for
      Compute $\mathcal{L}^{C}=\sum_{l}\mathcal{L}^{C}_{l}$;
       Compute output loss $\mathcal{L}^{O}$;
       Compute $L^{gen}=\mathcal{L}^{C}+\mathcal{L}^{O}$;
       Update $X^{*}_{\mathcal{B}}$ via backpropagation of $L^{gen}$;
end for
Algorithm 1 OuroMamba-Gen
@@ 5.1 Stage 1: OuroMamba-Gen
Following the typical DFQ setup Ramachandran et al. (2024a); Li et al. (2023b), OuroMamba-Gen requires an input batch of $\mathcal{B}$ random Gaussian noise images $X_{\mathcal{B}}$, and corresponding randomly generated task-specific targets $T_{G_{\mathcal{B}}}$ (detailed in section 6). $X_{\mathcal{B}}$ is fed to the VMM to minimize the generation loss $L^{gen}$ (described below), updating $X_{\mathcal{B}}$ via backpropagation for $G$ iterations. Upon convergence, the updated $X^{*}_{\mathcal{B}}$ is the generated synthetic data employed for calibration. The algorithm of OuroMamba-Gen pipeline is given in algorithm 1.
Building on the observations in section 4.1, OuroMamba-Gen first constructs an implicit attention representation that possesses enhanced global, spatial interactions for improved synthetic data generation.
@@
where, $w_{k}$ is a weighting factor that modulates the contribution of each neighboring state. Through our experiments and insights from prior work Han et al. (2025), we observe that $\Delta(t)$ exhibits higher magnitude responses in informative regions, such as foreground, while suppressing less relevant regions (Figure 1(b)). This property makes $\Delta(t)$ a natural choice for weighting the linear aggregation in Equation 7, enabling adaptive feature aggregation. To leverage this, we determine the weighting factor for all states in $\mathcal{N}(\tau)$ by performing a mean reduction along the channel dimension ($E$) of each $\Delta(\tau)$, as given by, $w_{k}=\textit{mean}_{E}(\Delta(k))$ for $k\in\mathcal{N}(\tau)$.
@@
Enhanced Implicit Attention. By substituting $h(t)$ with $h_{p}(t)$ in section 3.2, we obtain an enhanced representation of the N-dimensional implicit attention matrix ($\alpha$), denoted as $\alpha_{p}$. Interestingly, empirical validations show its ability to effectively separate the foreground from background (refer to Figure 1(d)), making it a potentially more informative representation compared to $\alpha$.
@@
Generation Loss ($\mathcal{L}^{gen}$). Ramachandran et al. (2024a) applied patch-level contrastive loss to self-attention in generating synthetic data for ViT. Inspired by this, we apply the contrastive loss (Equation 6) on $\alpha_{p}$. We specifically choose the patch-level contrastive loss as it has been proven by prior work to generate high quality synthetic data through capturing semantic relations. For a VMM layer $l$, each $\alpha^{l}_{p}[i,j]$ serves as the anchor patch i.e, $\lambda^{*}_{l,(i,j)}=\alpha^{l}_{p}[i,j]$ with positive and negative patches selected from spatially adjacent $\alpha^{l}_{p}[x,y],(x,y)\in\mathcal{N}(i,j)$. Similar to Ramachandran et al. (2024a), we leverage the cosine similarity metric to identify positive and negative patches in $\mathcal{N}(i,j)$, where we select the top-$n$ patches in $\mathcal{N}(i,j)$ with highest similarity as positive and rest as negative, where $n=\lfloor|\mathcal{N}(i,j)|/2\rfloor$. We then compute $\mathcal{L}^{C}_{l,(i,j)}$ for all anchor patches in $\alpha^{l}_{p}$ across all layers to get the final contrastive learning objective $\mathcal{L}^{C}$.
@@
Furthermore, to direct the synthetic data generation process towards task-specific goals Li et al. (2023b); Ramachandran et al. (2024a), we employ an additional output loss $\mathcal{L}^{O}$, which is the mean absolute error (MAE) between the predicted model output with the task specific targets. The final data generation loss is thus $L^{gen}=\mathcal{L}^{C}+\mathcal{L}^{O}$.
Input : Activation $X(t)\in\mathbb{R}^{N\times E}$, Static scale $S^{I}(t)$, Threshold $\theta$, Refresh rate $n_{\texttt{refresh}}$, Outlier list $O_{\texttt{list}}$, Inlier and outlier bit-precision $b_{a}^{I},b_{a}^{O}$
Output : Quantized activation $X_{q}(t)$, Updated outlier list $O_{\texttt{list}}$
if t % $n_{\texttt{refresh}}==0$ then
      $O_{\texttt{list}}$ = $\{\phi\}$
end if
$S^{D}(t)$ = ComputeScale($X(t)[:,c]\ \forall\ c\notin O_{\text{list}}$)
if $S^{D}(t)>S^{I}(t)$  then
      for each channel $c$ in $X(t)$ not in $O_{\text{list}}$ do
            if $max(|X(t)[:,c]|)\geq\theta$ then
                  $O_{\texttt{list}}$ = $O_{\texttt{list}}$cup $\{c\}$
             end if
       end for
end if
$I(t),O(t)$ = Separate($X(t)$, $O_{\texttt{list}}$)
$I_{q}(t)$ = InlierQuant($I(t),S^{I}(t),b_{a}^{I}$)
$O_{q}(t)$ = OutlierQuant($O(t),b_{a}^{O}$)
$X_{q}(t)=$Merge($I_{q}(t),O_{q}(t)$)
return $X_{q}(t)$, $O_{\texttt{list}}$
Algorithm 2 OuroMamba-Quant
@@ 5.2 Stage 2: OuroMamba-Quant
Building on the insights from section 4.2, OuroMamba-Quant is designed to accommodate the dynamic nature of VMMs. OuroMamba-Quant’s algorithm for activation quantization per-time-step is given in algorithm 2.
@@
Mixed Precision Quantization. To prevent significant accuracy degradation, we keep dynamically identified outliers at higher precision as compared to the inliers. Therefore, at every time-step each outlier channel having dedicated scale factor, are symmetric quantized to $b_{a}^{O}$-bits via per-channel quantization. On the other hand, inliers having a shared scale factor $S^{I}(t)$ over the tensor, are symmetric group quantized to $b_{a}^{I}$-bits ($b_{a}^{I}<b_{a}^{O}$) following Equation 5.
Weight Quantization. We apply per-channel symmetric group quantization for the weights. In specific, we follow Equation 5 to apply $b_{w}$-bit per-channel static quantization, where each channel $c$ has a shared scale factor $S^{W}_{c}$.
@@ 6.1 Experimental Setup
Models and Datasets. We evaluate OuroMamba across a range of VMM and hybrid Transformer-Mamba models, covering fundamental CV tasks including classification, detection, segmentation, and image generation tasks, as detailed below. Please see appendix on details of extension of OuroMamba-Quant to transformer layers.
Image Classification. We employ ImageNet-1K Deng et al. (2009) with 50K tests for ViM-T/S/B Zhu et al. (2024), VMamba-T/B Liu et al. (2025), LVMamba-S Huang et al. (2024) and hybrid MambaVision-T/B Hatamizadeh & Kautz (2025).
Object Detection. We use the COCO 2017 dataset Lin et al. (2015) having 20K test data. Following Ramachandran et al. (2024a); Li et al. (2023b), we use the Mask R-CNN He et al. (2017) framework with VMamba-S Liu et al. (2025) and MambaVision-T Hatamizadeh & Kautz (2025) as the backbones.
Semantic Segmentation. We avail the ADE20K dataset Zhou et al. (2019) with 3K test data encompassing 150 categories with VMamba-S Liu et al. (2025) and MambaVision-T Hatamizadeh & Kautz (2025) as the backbones. We employ the UperNet framework Xiao et al. (2018).
Image Generation. We evaluate generation performance using two high-resolution datasets FacesHQ Karras et al. (2019) and LandscapesHQ Skorokhodov et al. (2021) on Zigma, a zigzag Mamba backbone based diffusion model Hu et al. (2024).
Baselines for Comparison. We compare OuroMamba with state-of-the-art (SoTA) VMM PTQ techniques, PTQ4VM Cho et al. (2025) and QMamba Li et al. (2025b). Additionally, we compare with ViT PTQ methods, BREC-Q Li et al. (2021) and DopQ-ViT Yang et al. (2024), applying them to VMMs for a comprehensive evaluation.
@@
Task-Specific Synthetic Data Generation. For calibration, we use $\mathcal{B}=128$ synthetic samples. Following Ramachandran et al. (2024a), for image classification on the ImageNet-1K, we create $T_{G_{\mathcal{B}}}\in\mathcal{R}^{\mathcal{B}\times 1000}$, where the class-wise probabilities are randomly determined and assigned. Similarly, the target for object detection is $T_{G_{\mathcal{B}}}\in\mathcal{R}^{\mathcal{B}\times bb\times 5}$ where $bb$ is the number of bounding boxes in the image that is randomly selected from the integer set $[1,3]$. $T_{G_{\mathcal{B}}}[\mathcal{B},:,0]$ corresponds to the bounding box category and $T_{G_{\mathcal{B}}}[\mathcal{B},:,1:4]$ is the bounding box coordinates $x,y,w,h$ Huang et al. (2018). For segmentation, the target is a pixel-wise classification map of the same size as $X_{\mathcal{B}}$.
Quantization Setting Notation. Following prior work Lin et al. (2024); Chen et al. (2024c), the notation for a mixed-precision scheme such as OuroMamba is, W($b_{w}$)A($b_{a}^{I}$)O($b_{a}^{O}$). Since, in our evaluations we fix $b_{a}^{O}=8$, we omit $O8$ from the results for brevity.
@@
  Vim-S FP Baseline - - 32/32 81.60 32/32 81.60
BRECQ Li et al. (2021) R 1024 4/8 57.32 4/4 22.46
DopQ-ViT Yang et al. (2024) R 1024 4/8 68.96 4/4 61.58
PTQ4VM Cho et al. (2025) R 256 4/8 74.37 4/4 69.60
QMamba Li et al. (2025b) R 1024 4/8 74.12 4/4 33.64
@@
  Vim-B Baseline - - 32/32 81.90 32/32 81.90
BRECQ Li et al. (2021) R 1024 4/8 61.52 4/4 27.34
DopQ-ViT Yang et al. (2024) R 1024 4/8 68.47 4/4 62.33
PTQ4VM Cho et al. (2025) R 256 4/8 71.02 4/4 55.60
QMamba Li et al. (2025b) R 1024 4/8 75.46 4/4 66.73
@@
      VMamba-B FP Baseline - - 32/32 83.90 32/32 83.90
BRECQ Li et al. (2021) R 1024 4/8 62.34 4/4 25.65
DopQ-ViT Yang et al. (2024) R 1024 4/8 68.12 4/4 61.48
PTQ4VM Cho et al. (2025) R 256 4/8 78.95 4/4 75.67
QMamba Li et al. (2025b) R 1024 4/8 76.12 4/4 59.35
@@
      LVMamba-S FP Baseline - - 32/32 83.70 32/32 83.70
BRECQ Li et al. (2021) R 1024 4/8 55.34 4/4 11.62
DopQ-ViT Yang et al. (2024) R 1024 4/8 61.37 4/4 54.96
PTQ4VM Cho et al. (2025) R 256 4/8 81.46 4/4 78.21
QMamba Li et al. (2025b) R 1024 4/8 75.19 4/4 64.24
@@
      Hybrid Model MambaVision-B FP Baseline - - 32/32 84.20 32/32 84.20
BRECQ Li et al. (2021) R 1024 4/8 64.59 4/4 47.38
DopQ-ViT Yang et al. (2024) R 1024 4/8 70.68 4/4 61.43
PTQ4VM Cho et al. (2025) R 256 4/8 76.51 4/4 71.27
QMamba Li et al. (2025b) R 1024 4/8 76.08 4/4 68.59
@@ 6.2 Analysis of Generated Synthetic Data
In Figure 5, we visualize the synthetic samples generated by OuroMamba-Gen for Vim-B Zhu et al. (2024) for image classification (refer appendix for additional visualizations). Evidently, OuroMamba-Gen generates, realistic and clear class-specific foreground objects in contextually suitable background, showcasing a sophisticated understanding of semantic relationships between patches. This can be attributed to proposed enhanced impliict attention representation that is able to capture complex spatial relationships. This is in stark contrast to the naive synthetic samples in fig. 4(b), which are noisy and lack semantic structure.
@@
  VMamba-S MambaVision-T
@@
  BREC-Q Li et al. (2021) R 4/8 35.6 33.7 4/8 39.9 40.1
DopQ-ViT Yang et al. (2024) R 4/8 39.2 39.1 4/8 40.8 40.5
PTQ4VM Cho et al. (2025) R 4/8 48.4 42.0 4/8 46.2 40.9
QMamba Li et al. (2025b) R 4/8 47.2 41.3 4/8 45.3 39.9
@@
  BREC-Q Li et al. (2021) R 4/4 27.3 25.4 4/4 29.6 25.1
DopQ-ViT Yang et al. (2024) R 4/4 31.5 31.1 4/4 30.9 29.7
PTQ4VM Cho et al. (2025) R 4/4 45.6 41.3 4/4 43.8 39.4
QMamba Li et al. (2025b) R 4/4 43.7 39.1 4/4 42.1 38.6
@@ 6.3 Quantization Results for Classification
  VMamba-S MambaVision-T
@@
As shown in Table 1, existing ViT PTQ techniques Yang et al. (2024); Li et al. (2021) suffer from significant accuracy degradations–particularly at W4A4–of up to $72\%$ on VMM models and up to $37\%$ on hybrid models. Clearly demonstrating the limitations of ViT PTQ methods in handling the dynamic outlier characteristics of VMMs. On the other hand, baseline VMM PTQ methods PTQ4VM Cho et al. (2025), QMamba Li et al. (2025b) achieve acceptable performance at W4A8. However, at W4A4, they face large accuracy drops of up to $47\%$, due to static temporal grouping, per-token static scale factors. Furthermore, PTQ4VM’s migration of outliers from activations to weights, complicates weight quantization, contributing to its accuracy degradation. In contrast, OuroMamba consistently outperforms baselines across different models and W/A settings, achieving near-lossless performance, with an accuracy drop of $\leq 1.8\%$ at W4A8. Similarly, at W4A4 OuroMamba yields an accuracy improvement of $\mathbf{7.84}\%$ over PTQ4VM and $\mathbf{19.40}\%$ over QMamba on average, at W4A4. It is important to note that, OuroMamba’s superior performance is achieved in a data-free scenario requiring only 128 synthetic data samples, compared to PTQ4VM and QMamba requiring 256 and 1024 real calibration data.

### v2

@@ 4.2 Quantization
It has been well established by prior work [dong2023packqvit, jiang2025adfq, yang2024dopq] that outlier channels in ViT activations follow a fixed pattern, allowing anticipation in advance during calibration. However, unlike ViTs, VMMs exhibit dynamic variations in activation outlier channels across time-steps. As shown in Fig. 3, a comparison of activations $\bar{A},\bar{B}$ at time-steps 0 and 130 reveals that outlier channels (in red) differ significantly between time-steps. This temporal variance in outlier positions suggests that a static, pre-determined calibration approach, as used in ViTs, is insufficient for VMM quantization. Therefore, VMM necessitates dynamic identification of outlier channels at inference to reduce quantization error.
@@ 5.1 Stage 1: OuroMamba-Gen
Following the typical DFQ setup [ramachandran2024clamp, li2023psaq], OuroMamba-Gen requires an input batch of $\mathcal{B}$ random Gaussian noise images $X_{\mathcal{B}}$, and corresponding randomly generated task-specific targets $T_{G_{\mathcal{B}}}$ (detailed in Sec. 6). $X_{\mathcal{B}}$ is fed to the VMM to minimize the generation loss $L^{gen}$ (described below), updating $X_{\mathcal{B}}$ via backpropagation for $G$ iterations. Upon convergence, the updated $X^{*}_{\mathcal{B}}$ is the generated synthetic data employed for calibration. The algorithm of OuroMamba-Gen pipeline is given in supplementary.
Building on the observations in Sec. 4.1, OuroMamba-Gen first constructs an implicit attention representation that possesses enhanced global, spatial interactions for improved synthetic data generation.
@@
where, $w_{k}$ is a weighting factor that modulates the contribution of each neighboring state. Through our experiments and insights from prior work [han2025demystify], we observe that $\Delta(t)$ exhibits higher magnitude responses in informative regions, such as foreground, while suppressing less relevant regions (LABEL:fig:teaser(b)). This property makes $\Delta(t)$ a natural choice for weighting the linear aggregation in Eq. 7, enabling adaptive feature aggregation. To leverage this, we determine the weighting factor for all states in $\mathcal{N}(\tau)$ by performing a mean reduction along the channel dimension ($E$) of each $\Delta(\tau)$, as given by, $w_{k}=\textit{mean}_{E}(\Delta(k))$ for $k\in\mathcal{N}(\tau)$.
@@
Enhanced Implicit Attention. By substituting $h(t)$ with $h_{p}(t)$ in Sec. 3.2, we obtain an enhanced representation of the N-dimensional implicit attention matrix ($\alpha$), denoted as $\alpha_{p}$. Interestingly, empirical validations show its ability to effectively separate the foreground from background (refer to LABEL:fig:teaser(d)), making it a potentially more informative representation compared to $\alpha$.
@@
Generation Loss ($\mathcal{L}^{gen}$). [ramachandran2024clamp] applied patch-level contrastive loss to self-attention in generating synthetic data for ViT. Inspired by this, we apply the contrastive loss (Eq. 6) on $\alpha_{p}$. We specifically choose the patch-level contrastive loss as it has been proven by prior work to generate high quality synthetic data through capturing semantic relations. For a VMM layer $l$, each $\alpha^{l}_{p}[i,j]$ serves as the anchor patch i.e, $\lambda^{*}_{l,(i,j)}=\alpha^{l}_{p}[i,j]$ with positive and negative patches selected from spatially adjacent $\alpha^{l}_{p}[x,y],(x,y)\in\mathcal{N}(i,j)$. Similar to [ramachandran2024clamp], we leverage the cosine similarity metric to identify positive and negative patches in $\mathcal{N}(i,j)$, where we select the top-$n$ patches in $\mathcal{N}(i,j)$ with highest similarity as positive and rest as negative, where $n=\lfloor|\mathcal{N}(i,j)|/2\rfloor$. We then compute $\mathcal{L}^{C}_{l,(i,j)}$ for all anchor patches in $\alpha^{l}_{p}$ across all layers to get the final contrastive learning objective $\mathcal{L}^{C}$.
@@
Furthermore, to direct the synthetic data generation process towards task-specific goals [li2023psaq, ramachandran2024clamp], we employ an additional output loss $\mathcal{L}^{O}$, which is the mean absolute error (MAE) between the predicted output with the task specific targets. The final generation loss is thus $L^{gen}=\mathcal{L}^{C}+\mathcal{L}^{O}$.
@@ 5.2 Stage 2: OuroMamba-Quant
Building on the insights from Sec. 4.2, OuroMamba-Quant is designed to accommodate the dynamic nature of VMMs. OuroMamba-Quant’s algorithm for activation quantization per-time-step is given in supplementary.
@@
Mixed Precision Quantization. To prevent significant accuracy degradation, we keep dynamically identified outliers at higher precision as compared to the inliers. Therefore, at every time-step each outlier channel having dedicated scale factor, are symmetric quantized to $b_{a}^{O}$-bits via per-channel quantization. On the other hand, inliers having a shared scale factor $S^{I}(t)$ over the tensor, are symmetric group quantized to $b_{a}^{I}$-bits ($b_{a}^{I}<b_{a}^{O}$) following Eq. 5.
Weight Quantization. We apply per-channel symmetric group quantization for the weights. In specific, we follow Eq. 5 to apply $b_{w}$-bit per-channel static quantization, where each channel $c$ has a shared scale factor $S^{W}_{c}$.
@@ 6.1 Experimental Setup
Models and Datasets. We evaluate OuroMamba across a range of VMM and hybrid Transformer-Mamba models, covering fundamental CV tasks including classification, detection, segmentation, and image generation tasks, as detailed below. Please see supplementary on details of extension of OuroMamba-Quant to transformer layers.
Image Classification. We employ ImageNet-1K [deng2009imagenet] with 50K tests for ViM-T/S/B [zhu2024vision], VMamba-T/B [liu2025vmamba], LVMamba-S [huang2024localmamba] and hybrid MambaVision-T/B [hatamizadeh2024mambavision].
Object Detection. We use the COCO 2017 dataset [lin2015coco] having 20K test data. Following [ramachandran2024clamp, li2023psaq], we use the Mask R-CNN [he2017mask] framework with VMamba-S [liu2025vmamba] and MambaVision-T [hatamizadeh2024mambavision] as the backbones.
Semantic Segmentation. We avail the ADE20K dataset [zhou2019semantic] with 3K test data encompassing 150 categories with VMamba-S [liu2025vmamba] and MambaVision-T [hatamizadeh2024mambavision] as the backbones. We employ the UperNet framework [xiao2018unified].
Image Generation. We evaluate performance using two datasets FacesHQ [FFHQ_NVIDIA] and LandscapesHQ [2021landscape] on Zigma, a zigzag Mamba backbone based diffusion model [hu2024zigma].
Baselines for Comparison. We compare OuroMamba with state-of-the-art (SoTA) VMM PTQ techniques, PTQ4VM [cho2024ptq4vm] and QMamba [li2025qmamba]. Additionally, we compare with ViT PTQ methods, BREC-Q [li2021brecq] and DopQ-ViT [yang2024dopq], applying them to VMMs for a comprehensive evaluation.
@@
Task-Specific Synthetic Data Generation. For calibration, we use $\mathcal{B}=128$ synthetic samples. Following [ramachandran2024clamp], for image classification on the ImageNet-1K, we create $T_{G_{\mathcal{B}}}\in\mathcal{R}^{\mathcal{B}\times 1000}$, where the class-wise probabilities are randomly determined and assigned. Similarly, the target for object detection is $T_{G_{\mathcal{B}}}\in\mathcal{R}^{\mathcal{B}\times bb\times 5}$ where $bb$ is the number of bounding boxes in the image that is randomly selected from the integer set $[1,3]$. $T_{G_{\mathcal{B}}}[\mathcal{B},:,0]$ corresponds to the bounding box category and $T_{G_{\mathcal{B}}}[\mathcal{B},:,1:4]$ is the bounding box coordinates $x,y,w,h$ [huang2018introduction]. For segmentation, the target is a pixel-wise classification map of the same size as $X_{\mathcal{B}}$.
Quantization Setting Notation. Following prior work [lin2024qserve, 2024mixq, ramachandran2024algorithm], the notation for a mixed-precision scheme such as OuroMamba is, W($b_{w}$)A($b_{a}^{I}$)O($b_{a}^{O}$). Since, in our evaluations we fix $b_{a}^{O}=8$, we omit $O8$ from the results for brevity.
@@
  Vim-S [zhu2024vision] FP Baseline - - 32/32 81.60 32/32 81.60
BRECQ [li2021brecq] R 1024 4/8 57.32 4/4 22.46
DopQ-ViT [yang2024dopq] R 1024 4/8 68.96 4/4 61.58
PTQ4VM [cho2024ptq4vm] R 256 4/8 74.37 4/4 69.60
QMamba [li2025qmamba] R 1024 4/8 74.12 4/4 33.64
@@
  Vim-B [zhu2024vision] Baseline - - 32/32 81.90 32/32 81.90
BRECQ [li2021brecq] R 1024 4/8 61.52 4/4 27.34
DopQ-ViT [yang2024dopq] R 1024 4/8 68.47 4/4 62.33
PTQ4VM [cho2024ptq4vm] R 256 4/8 71.02 4/4 55.60
QMamba [li2025qmamba] R 1024 4/8 75.46 4/4 66.73
@@
      VMamba-B [liu2025vmamba] FP Baseline - - 32/32 83.90 32/32 83.90
BRECQ [li2021brecq] R 1024 4/8 62.34 4/4 25.65
DopQ-ViT [yang2024dopq] R 1024 4/8 68.12 4/4 61.48
PTQ4VM [cho2024ptq4vm] R 256 4/8 78.95 4/4 75.67
QMamba [li2025qmamba] R 1024 4/8 76.12 4/4 59.35
@@
      LVMamba-S [huang2024localmamba] FP Baseline - - 32/32 83.70 32/32 83.70
BRECQ [li2021brecq] R 1024 4/8 55.34 4/4 11.62
DopQ-ViT [yang2024dopq] R 1024 4/8 61.37 4/4 54.96
PTQ4VM [cho2024ptq4vm] R 256 4/8 81.46 4/4 78.21
QMamba [li2025qmamba] R 1024 4/8 75.19 4/4 64.24
@@
      Hybrid Model MambaVision-B [hatamizadeh2024mambavision] FP Baseline - - 32/32 84.20 32/32 84.20
BRECQ [li2021brecq] R 1024 4/8 64.59 4/4 47.38
DopQ-ViT [yang2024dopq] R 1024 4/8 70.68 4/4 61.43
PTQ4VM [cho2024ptq4vm] R 256 4/8 76.51 4/4 71.27
QMamba [li2025qmamba] R 1024 4/8 76.08 4/4 68.59
@@ 6.2 Analysis of Generated Synthetic Data
In Fig. 4, we visualize the synthetic samples generated by OuroMamba-Gen for Vim-B [zhu2024vision] for image classification (refer supplementary for additional visualizations). Evidently, OuroMamba-Gen generates, realistic and clear class-specific foreground objects in contextually suitable background, showcasing a sophisticated understanding of semantic relationships between patches. This can be attributed to proposed enhanced impliict attention representation that is able to capture complex spatial relationships. This is in stark contrast to the naive synthetic samples in Fig. 2(b), which are noisy and lack semantic structure.
@@
  VMamba-S [liu2025vmamba] MambaVision-T [hatamizadeh2024mambavision]
@@
  BREC-Q [li2021brecq] R 4/8 35.6 33.7 4/8 39.9 40.1
DopQ-ViT [yang2024dopq] R 4/8 39.2 39.1 4/8 40.8 40.5
PTQ4VM [cho2024ptq4vm] R 4/8 48.4 42.0 4/8 46.2 40.9
QMamba [li2025qmamba] R 4/8 47.2 41.3 4/8 45.3 39.9
@@
  BREC-Q [li2021brecq] R 4/4 27.3 25.4 4/4 29.6 25.1
DopQ-ViT [yang2024dopq] R 4/4 31.5 31.1 4/4 30.9 29.7
PTQ4VM [cho2024ptq4vm] R 4/4 45.6 41.3 4/4 43.8 39.4
QMamba [li2025qmamba] R 4/4 43.7 39.1 4/4 42.1 38.6
@@
  VMamba-S [liu2025vmamba] MambaVision-T [hatamizadeh2024mambavision]
@@
6.3 Quantization Results for Classification
As shown in Table 1, existing ViT PTQ techniques [yang2024dopq, li2021brecq] suffer from significant accuracy degradations–particularly at W4A4–of up to $72\%$ on VMM models and up to $37\%$ on hybrid models. Clearly demonstrating the limitations of ViT PTQ methods in handling the dynamic outlier characteristics of VMMs. On the other hand, baseline VMM PTQ methods PTQ4VM [cho2024ptq4vm], QMamba [li2025qmamba] achieve acceptable performance at W4A8. However, at W4A4, they face large accuracy drops of up to $47\%$, due to static temporal grouping, per-token static scale factors. Furthermore, PTQ4VM’s migration of outliers from activations to weights, complicates weight quantization, contributing to its accuracy degradation. In contrast, OuroMamba consistently outperforms baselines across different models and W/A settings, achieving near-lossless performance, with an accuracy drop of $\leq 1.8\%$ at W4A8. Similarly, at W4A4 OuroMamba yields an accuracy improvement of $\mathbf{7.84}\%$ over PTQ4VM and $\mathbf{19.40}\%$ over QMamba on average, at W4A4. It is important to note that, OuroMamba’s superior performance is achieved in a data-free scenario requiring only 128 synthetic data samples, compared to PTQ4VM and QMamba requiring 256 and 1024 real calibration data.